OpenStudy (anonymous):

What is the solution of the system? Use substitution. y=x+6y=2x A.(6,12) B.(-12,-6) C.(-6,-12) D.(2,4)

OpenStudy (anonymous):

y=x+6 x=2x By substitution, i get 2x=x+6 2x - x = x - x + 6 x=6 and get (6,12) as the solution

OpenStudy (anonymous):

That's right thank you.

OpenStudy (anonymous):

no problem ^_^